OpenStudy (yrelhan4):

|z-4|=re(z) maximim +arg of complex satisfying this??

OpenStudy (yrelhan4):

oh, maximum positive argument of z satisfying |z-4|=re(z)? i am sorry. messed up a little

OpenStudy (anonymous):

mod((x+iy)-4)=x mod((x-4)+iy))=x sqrt((x-4)^2+y^2)=x

OpenStudy (yrelhan4):

i get 16-8x+y^2.. how do i proceed?

OpenStudy (yrelhan4):

=0 *

OpenStudy (yrelhan4):

@nitz ??

OpenStudy (anonymous):

sorry that i was thinking could be a way to solve problem but i was wrong....................now i have found a way to solve it .................but dont know how to draw figure here??????????

OpenStudy (yrelhan4):

you can draw it by clicking draw.. right below the box yu type in..

OpenStudy (yrelhan4):

@nitz can you help me please?

OpenStudy (yrelhan4):

options are 30 60,45,90..

OpenStudy (experimentx):